# Beyond Classical Thermodynamics: Insight into the Microstructures of Black Holes with Quantum-Triggered Violations

### v1

@@ I Introduction
Since the pioneering work that established the four laws of black hole thermodynamics Bardeen:1973gs , the field of black hole physics has evolved significantly. The identification of surface gravity and area as black hole temperature Hawking:1975vcx and entropy Bekenstein:1973ur , respectively, has opened up new avenues of research in black hole thermodynamics. One of the most vibrant research areas in this domain centers on exploring phase transitions in black holes, particularly within the fascinating realm of black hole chemistry in the extended phase space framework, where the cosmological constant is viewed as a thermodynamic pressure variable, i.e., $P=-\Lambda/8\pi$ Kastor:2009wy ; Dolan:2010ha ; Dolan:2011 . With such an interpretation, in Ref.Kubiznak:2012wp the authors has demonstrated that four-dimensional charged AdS (Anti-de Sitter) black holes exhibit both first-order (discontinuous) and second-order (continuous) phase transitions between small and large black hole states within this extended phase space. This transition mirrors the Van der Waals gas/liquid phase transition, resulting in identical critical exponents for both systems, and hence obey the thermodynamic scaling laws as described by mean field theory. The study of critical behavior in black holes within this context, known as “$P$-$V$ criticality” has been extensively explored in the literature Gunasekaran:2012dq ; Wei:2012ui ; Cai:2013qga ; Dehghani:2014caa ; Hendi:2012um (see also Xu:2015rfa ; Cai:2014znn ; Majhi:2016txt ; Cheng:2016bpx ; Dehyadegari:2017hvd ; Kumara:2019xgt ; Li:2020xkh ; Hu:2018qsy ; Bhattacharya:2017hfj ; Bhattacharya:2017nru ; Hendi:2017fxp for more related works and reviews Altamirano:2014tva ; Kubiznak:2016qmn ). Various intriguing phenomena have been observed in the extended phase space of black holes, including multiple critical points Tavakoli:2022kmo , triple critical points Altamirano:2013uqa , reentrant phase transitions Altamirano:2013ane , and superfluid-like phase transitions Hennigar:2016xwd .
@@
where the function $f(r)$ is given by [^1] In this scenario, the branch with the negative sign (``"``"``-"` ` - ") pertains to an AdS spacetime, whereas the branch with the positive sign (``+"``"``+"` ` + ") corresponds to a de Sitter spacetime. However, in this paper, we are solely interested in the AdS scenario and its thermodynamic properties.
@@
considering only type A anomaly with $\beta=0$. Here, $M$ and $Q$ are integration constants, while $l$ is linked to the cosmological constant via $\Lambda=-{3}/{l^{2}}$. In the limit of large $r$, where quantum corrections are negligible, this solution approximates to the Reissner-Nordström-Anti-de Sitter (RN-AdS) black hole. This black hole solution, which incorporates the effects of quantum conformal anomaly, exhibits interesting thermodynamic properties. By examining the thermodynamic properties of this black hole solution, a deeper insight can be gained into the black hole’s microscopic structure and the fundamental degrees of freedom that govern its behavior.
Interestingly, a recent studies has unveiled a fascinating discovery: in black hole systems with quantum conformal anomaly, the scaling laws can be violated under specific conditions Hu:2024ldp . When $\alpha_{\rm c}\leqslant 0$, the behavior of the black hole system aligns with mean field theory. In this case, the scaling laws, which describe the relationships between various physical quantities in the system, are upheld and not violated. This indicates that the presence of quantum conformal anomaly does not significantly alter the standard behavior expected from mean field theory under these conditions. Specifically, when $\alpha_{\rm c}={Q^{2}}/{8}$, an intriguing breakdown of scaling laws occurs in the black hole system. This specific condition arises when the coefficient $a_{11}$ in the expansion of pressure $P(V,T)$ near the critical point is zero. Here, the central charge parameter $\alpha_{\rm c}$ and $U(1)$ charge $Q$ are related in such a way that triggers a unique phase structure in the black hole system. This phase structure differs significantly from that observed in van der Waals-like systems, which typically adhere to scaling laws. The violation of scaling laws in this scenario suggests that the quantum conformal anomaly plays a crucial role in modifying the behavior of the black hole system under these specific conditions.
@@
The structure of the paper is organized as follows. In Sec. II, we shall make a brief review on Ruppeiner thermodynamic geometry. In Sec.III, we shall study the behavior of thermodynamic curvature for the charged AdS black hole with quantum anomaly, where we obtain the coexisting volumes of the black hole during the $P$-$V$ phase transition. In Sec.IV, we shall obtain the coexisting volumes of the charged AdS black hole with quantum anomaly during the $P$-$V$ phase transition by considering the specific relation $\alpha_{c}={Q^{2}}/{8}$, corresponding to the case that the black hole system exhibits a breakdown of scaling laws. We also study the behavior of thermodynamic curvature of this black hole along coexisting volumes. In Sec.V, we end the work with conclusions and discussion. Throughout this paper, we adopt natural units such that $c=G=\hbar=1$.
@@ II Warm-up: Ruppeiner thermodynamic geometry
In this section, we will provide a concise overview of some key results of Ruppeiner thermodynamic geometry Ruppeiner:1995zz ; Ruppeiner:2012uc ; Ruppeiner:2013yca ; Wei:2019uqg as an introduction. While statistical mechanics allows us to derive macroscopic thermodynamic quantities from the microscopic constituents of an ordinary fluid system, our current knowledge necessitates adopting a thermodynamic geometric approach in reverse to understand the microscopic structure of a black hole. We start with Boltzmann’s entropy formula Ruppeiner:1995zz
@@
Now, consider a thermodynamic system divided into a small subsystem $S$ and its environment $E$. The total entropy can be decomposed as $S(x^{0},...,x^{\rm N})=S_{\rm S}(x^{0},...,x^{\rm N})+S_{\rm E}(x^{0},...,x^{\rm N})$ with $S_{\rm S}\ll S_{\rm E}\sim S$. Expanding the total entropy near its local maximum at $x^{\mu}=x^{\mu}_{0}$, one obtains
@@
where $S_{0}$ represents the local maximum of the entropy and “$|_{0}$” denotes evaluation at $|_{x^{\mu}=x^{\mu}_{0}}$. It is reasonable to assume that the fluctuating parameters are both conserved and additive, meaning $x^{\mu}_{\rm S}+x^{\mu}_{\rm E}=x^{\mu}_{\rm total}$, which remains constant. Consequently, the sum of the second and third terms equals zero (or equivalently, $\frac{\partial S_{\rm S}}{\partial x^{\mu}}|_{0}\Delta x^{\mu}_{\rm S}$ is the negative of $\frac{\partial S_{\rm E}}{\partial x^{\mu}}|_{0}\Delta x^{\mu}_{\rm E}$). Given that $S_{\rm E}$ is comparable in scale to the entire system, the final term in equation (8) is substantially smaller than the fourth term and can therefore be neglected. Therefore, one can approximate the entropy change as
@@ III.1 Coexistence Curves of the $P$-$V$ Phase Transition
The necessary condition for the occurrence of a $P$-$V$ phase transition in the system described by Eq.(13) is that the first and second derivatives of the pressure $P$ with respect to the volume $V$ at constant temperature $T$ must both be zero:
@@
where $K\equiv\sqrt{192\alpha_{\rm c}^{2}+9Q^{4}-96\alpha_{\rm c}Q^{2}}$ and $V_{\rm c}=4\pi r_{\rm hc}^{3}/3$, with $r_{\rm hc}$ being the critical horizon radius. To conveniently and explicitly display the diagrams, we first define the dimensionless reduced pressure, reduced temperature, and reduced volume as follows
$\widetilde{P}\equiv\frac{P}{P_{\rm c}}\,,\quad\quad\widetilde{V}\equiv\frac{V}{V_{\rm c}}\,,\quad\quad\widetilde{T}\equiv\frac{T}{T_{\rm c}}\,.$ (16)
@@
Figure 1: $\widetilde{P}$-$\widetilde{V}$ isotherms for the reduced equation of state of a charged AdS black hole with quantum conformal anomaly. The diagram shows ideal gas behavior, indicating a unique phase of the black hole for $T>T_{\rm c}$, and oscillatory behavior for $T<T_{\rm c}$, which suggests the existence of small and large black hole phases undergoing a first-order phase transition.
@@
$a_{11}$ $=$ $\frac{2(5K-3Q^{2}+24\alpha_{\rm c})}{27Q^{2}-66\alpha_{\rm c}}\,,\,\,a_{03}=\frac{K-15Q^{2}+40\alpha_{\rm c}}{27(9Q^{2}-22\alpha_{\rm c}}\,.\,\,$
@@
[…]
@@ III.2 Thermodynamic curvature via $P$-$V$ criticality
$R_{\rm N}=\frac{2\delta\beta}{(\beta+\delta)^{2}}\,,$ (23)
@@
$\beta$ $=$ $\delta+\widetilde{T}\widetilde{V}(K^{2}-9Q^{4})-12\alpha_{\rm c}(K-3Q^{2})(1+\widetilde{V}^{2/3})\widetilde{T}{\widetilde{V}}^{1/3}\,$
@@
Note that $R_{\rm N}$ explicitly depend on the black hole charge $Q$ and quantum anomaly. We see that for most of the parameter space the value of $R_{\rm N}$ is near zero. However, in proximity to the critical point defined by the equation
@@
$R_{\rm N}$ undergoes significant variations, diving to negative infinity along this curve. This behavior suggests a rapid transformation in the microstructure of the charged AdS black hole with quantum anomaly near the temperature $\tilde{T}_{\rm div}$. Furthermore, we have computed the curves where $R_{\rm N}$ changes its sign, given by
@@
Finally, we discuss the behaviors of thermodynamic curvature towards the critical point along the coexistence curve. For this purpose, we write $R_{\rm N}$ as a function of $\widetilde{T}$ by using Eqs.(20), (21) and (23), the series expansions of it at $\widetilde{T}$ along the coexistence curves have the following forms
$R_{\rm N}^{s}$ $=$ $-\frac{1}{8t^{2}}-\frac{\eta}{{[(8\alpha_{\rm c}-3Q^{2})t]}^{3/2}}+...O(t^{-1},...)\,,$ (26)
$R_{\rm N}^{l}$ $=$ $-\frac{1}{8t^{2}}+\frac{\eta}{{[(8\alpha_{\rm c}-3Q^{2})t]}^{3/2}}+...O(t^{-1},...)\,,$ (27)
@@
$\eta=\frac{\sqrt{K}(768\alpha_{\rm c}^{2}-80\alpha_{\rm c}K+31KQ^{2}+63Q^{4}-456\alpha_{\rm c}Q^{2})}{4\sqrt{2}(-8\alpha_{\rm c}+K+Q^{2})}$ (28)
@@
This analysis unveils a universal, dimensionless constant of $-1/8$, which aligns perfectly with the numerical findings for the van der Waals fluid, as reported in Wei:2019uqg . To visualize how $R_{\rm N}$ diverges, we present a graphical representation of the normalized thermodynamic curvature for both the small and large coexistence phases in Fig.3.
@@
Figure 3: The normalized thermodynamic curvature $R_{\rm N}$ of the charged AdS black hole with quantum anomaly for $\alpha_{\rm c}\leqslant 0$ along the coexistence small and large phases.
@@
Figure 3 shows that the normalized thermodynamic curvature $R_{\rm N}$ in both small and large phases diverges to negative infinity ($R_{\rm N}\rightarrow-\infty$) at the critical point. Moreover, $\left|R_{\rm N}\right|$ decreases as the temperature deviated from the critical temperature, becoming small at small $\widetilde{T}$. Notably, at the critical point, the normalized thermodynamic curvature $R_{\rm N}$ diverges to negative infinity in both the small and large phases. As the temperature deviates from the critical temperature, $\left|R_{\rm N}\right|$ decreases, becoming negligible at lower temperatures. In contrast to a van der Waals fluid system, where weak repulsive interactions dominate throughout the parameter space above the coexistence curve.

### v2

# Insight into the Microstructures of Black Holes with Quantum-Triggered Violations

@@ I Introduction
Since the pioneering work that established the four laws of black hole thermodynamics Bardeen:1973gs , the field of black hole physics has evolved significantly. The identification of surface gravity and area as black hole temperature Hawking:1975vcx and entropy Bekenstein:1973ur , respectively, has opened up new avenues of research in black hole thermodynamics. One of the most vibrant research areas in this domain centers on exploring phase transitions in black holes, particularly within the fascinating realm of black hole chemistry in the extended phase space framework, where the cosmological constant is viewed as a thermodynamic pressure variable, $i.e.$, $P=-\Lambda/8\pi$ Kastor:2009wy ; Dolan:2010ha ; Dolan:2011 . With such an interpretation, in Ref.Kubiznak:2012wp the authors has demonstrated that four-dimensional charged AdS (Anti-de Sitter) black holes exhibit both first-order (discontinuous) and second-order (continuous) phase transitions between small and large black hole states within this extended phase space. This transition mirrors the van der Waals gas/liquid phase transition, resulting in identical critical exponents for both systems, and hence obey the thermodynamic scaling laws as described by mean field theory. The study of critical behavior in black holes within this context, known as “$P$-$V$ criticality” has been extensively explored in the literature Gunasekaran:2012dq ; Wei:2012ui ; Cai:2013qga ; Dehghani:2014caa ; Hendi:2012um (see also Xu:2015rfa ; Cai:2014znn ; Majhi:2016txt ; Cheng:2016bpx ; Dehyadegari:2017hvd ; Kumara:2019xgt ; Li:2020xkh ; Hu:2018qsy ; Bhattacharya:2017hfj ; Bhattacharya:2017nru ; Hendi:2017fxp ; Abdusattar:2025rdp for more related works and reviews Altamirano:2014tva ; Kubiznak:2016qmn ). Various intriguing phenomena have been observed in the extended phase space of black holes, including multiple critical points Tavakoli:2022kmo , triple critical points Altamirano:2013uqa , reentrant phase transitions Altamirano:2013ane , and superfluid-like phase transitions Hennigar:2016xwd .
@@
where the function $f(r)$ is given by [^1] In this scenario, the branch with the negative sign (``-") pertains to an AdS spacetime, whereas the branch with the positive sign (+``+") corresponds to a de Sitter spacetime. However, in this paper, we are solely interested in the AdS scenario and its thermodynamic properties.
@@
considering only type $A$ anomaly with $\beta=0$. Here, $M$ is the mass of the black hole, $Q$ can be interpreted as the $U(1)$ charge of the underlying conformal field theory, and $\alpha_{c}$ is the central charge, while $l$ is linked to the cosmological constant via $\Lambda=-{3}/{l^{2}}$. In the limit of large $r$, where quantum corrections are negligible, this solution approximates to the Reissner-Nordström-Anti-de Sitter (RN-AdS) black hole. This black hole solution, which incorporates the effects of quantum conformal anomaly, exhibits interesting thermodynamic properties. By examining the thermodynamic properties of this black hole solution, a deeper insight can be gained into the black hole’s microscopic structure and the fundamental degrees of freedom that govern its behavior.
Interestingly, a recent study has unveiled a fascinating discovery: in black hole systems with quantum conformal anomaly, the scaling laws can be violated under specific conditions Hu:2024ldp . When $\alpha_{\rm c}\leqslant 0$, the black hole system aligns with mean field theory: the scaling laws (describing relationships between the system’s physical quantities) remain intact, indicating quantum conformal anomaly does not significantly alter the standard behavior expected from mean field theory under these conditions. However, when $\alpha_{\rm c}={Q^{2}}/{8}$, a striking breakdown of scaling laws occurs in the black hole system. This specific condition arises when the coefficient $a_{11}$ in the expansion of the pressure $P(V,T)$ near the critical point is equal to zero. Here, the central charge parameter $\alpha_{\rm c}$ and $U(1)$ charge $Q$ exhibit a correlation that induces a unique phase structure in the black hole system. This phase structure differs significantly from that observed in van der Waals-like systems, which typically adhere to scaling laws. The violation of scaling laws in this scenario highlights that the quantum conformal anomaly plays a crucial role in modifying the behavior of the black hole system under these specific conditions. Notably, the critical phenomenon of non-standard critical exponents was first observed in another type of black hole system Frassino:2014pha , and subsequent work has proposed a new interpretation for the physical origin of such non-standard critical behaviors Ahmed:2022kyv , providing valuable insights for understanding the critical behavior of quantum-corrected black holes.
@@
The structure of the paper is organized as follows. In Sec. II, we shall make a brief review on Ruppeiner thermodynamic geometry. In Sec.III, we shall study the behavior of thermodynamic curvature for the charged AdS black hole with quantum anomaly with a central charge $\alpha_{\rm c}\leqslant 0$, where we obtain the coexisting volumes of the black hole during the $P$-$V$ phase transition. In Sec. IV, we investigate the $P$-$V$ criticality and the behavior of the thermodynamic curvature of black holes with a positive central charge ($0<\alpha_{c}<{Q^{2}}/{8}$). In Sec.V, we shall obtain the coexisting volumes of the charged AdS black hole with quantum anomaly during the $P$-$V$ phase transition by considering the specific relation $\alpha_{c}={Q^{2}}/{8}$, corresponding to the case that the black hole system exhibits a breakdown of scaling laws. We also study the behavior of thermodynamic curvature of this black hole along coexisting volumes. In Sec.VI, we end the work with conclusions and discussion. Throughout this paper, we adopt natural units such that $c=G=\hbar=1$.
@@ II Warm-up: Ruppeiner thermodynamic geometry
In this section, we will provide a concise overview of some key results of Ruppeiner thermodynamic geometry Ruppeiner:1995zz ; Ruppeiner:2012uc ; Ruppeiner:2013yca ; Wei:2019uqg ; Wei:2019yvs as an introduction. While statistical mechanics allows us to derive macroscopic thermodynamic quantities from the microscopic constituents of an ordinary fluid system, our current knowledge necessitates adopting a thermodynamic geometric approach in reverse to understand the microscopic structure of a black hole. We start with Boltzmann’s entropy formula Ruppeiner:1995zz
@@
Now, let us consider a thermodynamic system divided into a small subsystem $S$ and its environment $E$. The total entropy can be decomposed as $S(x^{0},...,x^{\rm N})=S_{\rm S}(x^{0},...,x^{\rm N})+S_{\rm E}(x^{0},...,x^{\rm N})$ with $S_{\rm S}\ll S_{\rm E}\sim S$. Expanding the total entropy near its local maximum at $x^{\mu}=x^{\mu}_{0}$, one obtains
@@
where $S_{0}$ represents the local maximum of the entropy and $|_{0}$” denotes evaluation at $|_{x^{\mu}=x^{\mu}_{0}}$. It is reasonable to assume that the fluctuating parameters are both conserved and additive, meaning $x^{\mu}_{\rm S}+x^{\mu}_{\rm E}=x^{\mu}_{\rm total}$, which remains constant. Consequently, the sum of the second and third terms equals zero (or equivalently, $\frac{\partial S_{\rm S}}{\partial x^{\mu}}|_{0}\Delta x^{\mu}_{\rm S}$ is the negative of $\frac{\partial S_{\rm E}}{\partial x^{\mu}}|_{0}\Delta x^{\mu}_{\rm E}$). Given that $S_{\rm E}$ is comparable in scale to the entire system, the final term in equation (8) is substantially smaller than the fourth term and can therefore be neglected. Therefore, one can approximate the entropy change as
@@ III.1 Coexistence curves of the $P$-$V$ phase transition
The necessary condition for the occurrence of a $P$-$V$ phase transition in the system described by Eq.(13) is that the first and second derivatives of the pressure $P$ with respect to the volume $V$ at constant temperature $T$ must both be zero
@@
where $K\equiv\sqrt{192\alpha_{\rm c}^{2}+9Q^{4}-96\alpha_{\rm c}Q^{2}}$ and $V_{\rm c}=4\pi r_{\rm hc}^{3}/3$, with $r_{\rm hc}$ being the critical horizon radius. To conveniently and explicitly display the diagrams, we first define the dimensionless reduced pressure, reduced volume, and reduced temperature as follows
$\widetilde{P}\equiv\frac{P}{P_{\rm c}}\,,\quad\quad\widetilde{V}\equiv\frac{V}{V_{\rm c}}\,,\quad\quad\widetilde{T}\equiv\frac{T}{T_{\rm c}}\,$, (16)
and rewrite (13) as
$\widetilde{P}$ $=$ $\frac{\widetilde{T}T_{\rm c}\big[\sqrt[3]{\pi}(6V_{\rm c}\widetilde{V})^{2/3}-16\pi\alpha_{\rm c}\big]}{6P_{\rm c}V_{\rm c}\widetilde{V}}$
$+\frac{2(6\pi)^{2/3}(Q^{2}-2\alpha_{\rm c})-3\sqrt[3]{6}(V_{\rm c}\widetilde{V})^{2/3}}{36\sqrt[3]{\pi}P_{\rm c}(V_{\rm c}\widetilde{V})^{4/3}}\,.$
@@
Figure 1: Isotherms in the $\widetilde{P}-$$\widetilde{V}$ plane for the reduced equation of state of a charged AdS black hole with quantum conformal anomaly, with $\widetilde{T}=T/T_{\rm c}$. The green, red, and dashed blue curves correspond to $\widetilde{T}=1.1$, $\widetilde{T}=1$, and $\widetilde{T}=0.9$, respectively, with $\alpha_{\rm c}=-0.5$ and $Q=1$ fixed throughout all plots.
The diagram reveals that for $T>T_{\rm c}$, the curves exhibit ideal gas behavior, indicating a unique phase of the black hole. For $T<T_{\rm c}$, the oscillatory behavior suggests the existence of small and large black hole phases, which undergo a first-order phase transition. The red curve (at $T=T_{\rm c}$) displays the expected inflection point characteristic of a critical point.
@@
$a_{11}$ $=$ $\frac{2(3Q^{2}-5K-24\alpha_{\rm c})}{27Q^{2}-66\alpha_{\rm c}}\,,\,\,a_{03}=\frac{K-15Q^{2}+40\alpha_{\rm c}}{27(9Q^{2}-22\alpha_{\rm c})}\,.\,\,$
@@
$\widetilde{P}(t,\omega_{s})=\widetilde{P}(t,\omega_{l})\,,~~~~\int_{\omega_{s}}^{\omega_{l}}\omega d\widetilde{P}=0\,.$ (20)
@@
Figure 2: Coexistence curves of small and large black hole phases in the $\widetilde{T}-$$\widetilde{V}$ diagram for charged AdS black holes with quantum anomaly. Here we set $\alpha_{\rm c}=-0.5,Q=1$.
@@ III.2 Thermodynamic curvature via $P$-$V$ criticality
$R_{\rm N}=\frac{2\delta\beta_{0}}{(\beta_{0}+\delta)^{2}}\,$, (24)
@@
$\beta_{0}$ $=$ $\delta+\widetilde{T}\widetilde{V}(K^{2}-9Q^{4})-12\alpha_{\rm c}(K-3Q^{2})(1+\widetilde{V}^{2/3})\widetilde{T}{\widetilde{V}}^{1/3}\,$
@@
Note that $R_{\rm N}$ explicitly depend on the black hole charge $Q$ and quantum anomaly. As $\widetilde{T}$ tends to 0, we observe that $R_{\rm N}$ converges to $1/2$ for both phases, reflecting a common asymptotic property in the low-temperature limit. However, in proximity to the critical point defined by the equation
@@
$R_{\rm N}$ undergoes significant variations, diving to negative infinity along this curve. This behavior suggests a rapid transformation in the microstructure of the charged AdS black hole with quantum anomaly near the temperature $\widetilde{T}_{\rm div}$. Furthermore, we have computed the curves where $R_{\rm N}$ changes its sign, given by
@@
Finally, we discuss the behavior of the thermodynamic curvature as it approaches the critical point along the coexistence curve. To this end, we first express the normalized scalar curvature $R_{\rm N}$ as a function of the reduced temperature $\widetilde{T}$ by substituting Eqs. (21), (22), and (24); we then derive its series expansions along the coexistence curve, which take the following forms
[…]
@@
$\eta_{0}=\frac{\sqrt{K}(768\alpha_{\rm c}^{2}-80\alpha_{\rm c}K+31KQ^{2}+63Q^{4}-456\alpha_{\rm c}Q^{2})}{4\sqrt{2}(-8\alpha_{\rm c}+K+Q^{2})}$ (29)
@@
This analysis unveils a universal, dimensionless constant of 1/8-1/8, which aligns perfectly with the numerical findings for the van der Waals fluid, as reported in Wei:2019uqg . To visualize how $R_{\rm N}$ diverges, we present a graphical representation of the normalized thermodynamic curvature for both the small and large coexistence phases in Fig.3.
@@
Figure 3: The normalized thermodynamic curvature $R_{\rm N}$ of the charged AdS black hole with quantum anomaly for $\alpha_{\rm c}<0$ along the coexistence small and large phases. Here we set $\alpha_{\rm c}=-0.5,Q=1$.
@@
Figure 3 shows that the normalized thermodynamic curvature $R_{\rm N}$ in both small and large phases diverges to negative infinity ($R_{\rm N}\rightarrow$-$\infty$) at the critical point. Specifically, for the small black hole phase (dashed red curve), $R_{\rm N}$ changes sign to positive as the temperature deviates from the critical temperature. For the large black hole phase (solid blue curve), $R_{\rm N}$ exhibits a distinct behavior, remaining nearly constant at small $\widetilde{T}$ when the temperature deviates from the critical temperature. This contrasts with a van der Waals fluid system, where weak repulsive interactions prevail throughout the parameter space above the coexistence curve.